OpenStudy (anonymous):

after exercising for a few minutes a person has a respiratory cycle for which the rate of air intake is v=1.75sin(pit/2) find the volume, in liters, of air inhaled during one cycle by integrating the function over the interval (0,2).

OpenStudy (dumbcow):

\[V = 1.75\int\limits_{0}^{2} \sin(\frac{\pi}{2} t) dt\] use substitution....... u = (pi/2)t du = pi/2 dt \[V = (1.75)(\frac{2}{\pi})\int\limits_{0}^{2} \sin (u) du\] \[V = (1.75)(\frac{2}{\pi})\left[ -\cos(\frac{\pi}{2} t) \right]\] then evaluate from 0 to 2